$SPAD/src/input evalex.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool evalex.output
)set message test on
)set message auto off
)clear all
-- Input for page PrefixEval
--S 1 of 3
cos(2)
--R
--R
--R   (1)  cos(2)
--R                                                    Type: Expression(Integer)
--E 1
-- Input for page PrefixEval
)clear all
--S 2 of 3
cos(2)
--R
--R
--R   (1)  cos(2)
--R                                                    Type: Expression(Integer)
--E 2
-- Input for page InfixEval
)clear all
--S 3 of 3
2 + 3.4
--R
--R
--R   (1)  5.4
--R                                                                  Type: Float
--E 3
)spool
)lisp (bye)

———-

References^{[1] nothing}

4